OpenStudy (acespeedfighter):

(06.07 MC) Which expressions are equivalent to 3(x + 3y + 2x − y )? (1 point) 3(3x + 2y) and 9x + 6y 3(3x + 4y) and 9x + 12y 3(x + 4y) and 3x + 12y 3(x + 2y) and 3x + 6y

OpenStudy (hayisforhorses):

WHY

OpenStudy (acespeedfighter):

why? what

OpenStudy (hayisforhorses):

I JUST ANSWERED THIS

OpenStudy (mathstudent55):

Can you simplify the given expression. Combine like terms inside the parentheses.

OpenStudy (hayisforhorses):

I'm so confused

OpenStudy (acespeedfighter):

hmm...

OpenStudy (fopkate):

The answer is the first one because you combine the like terms inside of the parenthesis first, and then you distribute the 3 to that equation.

OpenStudy (mathstudent55):

\(3(\color{red}{x} + \color{green}{3y} + \color{red}{2x} + \color{green}{− y} ) =\) \(3(\color{red}{x} + \color{red}{2x} + \color{green}{3y} + \color{green}{− y} ) \) What do you get inside the parentheses?

OpenStudy (acespeedfighter):

3 + 2 = 5

OpenStudy (acespeedfighter):

2 + 3 = 5

OpenStudy (mathstudent55):

What is x + 2x = ?

OpenStudy (mathstudent55):

What is 3y - y = ?

OpenStudy (acespeedfighter):

x + 2 = 3 then 3 - y = 2

OpenStudy (mathstudent55):

There is no x + 2. Also x + 2 is not 3 It is x + 2x which is equal to 3x.

OpenStudy (mathstudent55):

With the y's it is not 3 - y It is 3y - y = 2y

OpenStudy (mathstudent55):

That means that when we combine like terms inside the parentheses, we get 3(3x + 2y)

OpenStudy (mathstudent55):

Then we use the distributive property to get \(3(3x + 2y) = 9x + 6y\)

OpenStudy (acespeedfighter):

still dont get it